AB=(-6 a7 [12 is and B= 1 4 2 - 3 solving AB with your A matrix. Find A. Check your answer by
Answer #1

7 such that AB la b] his life as -) 49+12 b=-6. 2a +66=-3 -0. -59-156-29, a+36=-975: +12a+86--1815) -@. we get different valuwe get two different values of same equation which is not valid. Thabk you